A small motor starts on the first attempt 3/8 of the time, on the second attempt 1/4 of the time, and on the third attempt 1/4 of the time.
Let the random variable X be the number of tries it takes to start the motor. So, for instance, P(X=1) is 3/8.
Use the definition of a probability mass function to find the probability it takes more than 3 tries to start the motor.
P(X > 3) ?

Answer by ikleyn(52781)

.

P(X > 3)  is the complement of the sum  P(X=1) + P(X=2) + P(X=3).


Therefore,  P(X > 3) = 1 - 3/8 - 1/4 - 14 = 1/8.    ANSWER

Solved.